OpenStudy (anonymous):

Simplify 57 - 6^2 ÷ (2 + 1) • 4.

OpenStudy (anonymous):

(57-36)/3*4 21/12

OpenStudy (anonymous):

have you got it

OpenStudy (kropot72):

You need to perform the operations in the correct order. First remove brackets to get: \[57-6^{2}\div3\times 4\] Next get rid of the exponent to get: \[57-36\div3\times 4\] Then perform the division to get: \[57-12\times 4\] And multiply to get: \[57-48\] Finally subtract to get ?